OpenStudy (anonymous):

Ravi weighs 56.7kg. If he reduces his weight in the ratio 7:6, find his new weight.

OpenStudy (anonymous):

a boy weighs 56.7kg. If he reduces his weight in the ratio 7:6, find his new weight. : Follow the procedure and you will see why. : Let x = the new weight. write it as a ratio equation Old weight is to new weight; 56.7/x = 7/6 Cross multiply 7x = 56.7 * 6 7x = 340.2 x = 340.2/7 x = 48.6

OpenStudy (anonymous):

divide by seven and multiply by six... I think